# Carbon-Aware Computing for Data Centers with Probabilistic Performance Guarantees

### v2

@@ Abstract
Data centers are significant contributors to carbon emissions and can strain power systems due to their high electricity consumption. To mitigate this impact and to participate in demand response programs, cloud computing companies strive to balance and optimize operations across their global fleets by making strategic decisions about when and where to place compute jobs for execution. In this paper, we introduce a load shaping scheme which reacts to time-varying grid signals by leveraging both temporal and spatial flexibility of compute jobs to provide risk-aware management guidelines and job placement with provable performance guarantees based on distributionally robust optimization. Our approach divides the problem into two key components: (i) day-ahead planning, which generates an optimal scheduling strategy based on historical load data, and (ii) real-time job placement and (time) scheduling, which dynamically tracks the optimal strategy generated in (i). We validate our method in simulation using normalized load profiles from randomly selected Google clusters, incorporating time-varying grid signals. We can demonstrate significant reductions in carbon cost and peak power with our approach compared to myopic greedy policies, while maintaining computational efficiency and abiding to system constraints.
Index Terms:
Data-driven distributionally robust optimization, job scheduling, data center optimization, demand shift
@@ I Introduction
The literature covers various aspects of the problem but has key limitations. Specifically, most works (i) do not exploit temporal and spatial flexibility jointly [19, 20, 21, 22]; (ii) rely on either optimization-based planning approaches, which are not tractable in real-time operation [21, 23], or myopic greedy policies, which are suboptimal as they do not take into account future grid signals or compute load predictions [24, 16, 25]; and (iii) either assume exact knowledge of incoming compute jobs, which is unrealistic, or lack specifics on real-time job placement [21, 18].
@@
We propose a distributionally-robust day-ahead planning problem coupled with a real-time placement algorithm which is driven by data, offers robustness guarantees, and considers various degrees of flexibility of jobs. Specifically, our approach is unique in the following ways:
@@
    Fully data-driven: We directly approximate the probability distribution of compute loads from data within the planning problem.
@@
The rest of the paper is structured as follows. In Section II, we introduce the model and formulate the stochastic optimization problem. In Section III, we derive a distributionally-robust load schedule, approximating the ambiguity set from historical data. Section IV discusses real-time job placement. Section V presents simulation results and Section VI concludes the paper.
I-A Basic Notation
We denote the set of the first $K$ positive integers as $\mathbb{Z}_{K}:=\{1,\dots,K\}$, and the set of positive integers from $t$ to $T$, with $t<T$, as $\mathbb{Z}_{[t:T]}:=\{t,\dots,T\}$. The operator $[\,\cdot\,]_{+}:=\max\{\cdot,0\}$ is the projection on the positive orthant. Given some variables $v_{t,d}$, with $t\in\mathcal{T}$ and $d\in\mathcal{D}$, we denote the stacked vector as $v:=\text{col}(v_{t,d})_{t\in\mathcal{T},d\in\mathcal{D}}=[v_{1,1},\dots,v_{T,D}]^{\top}$. Table I summarizes all parameters, sets, and variables used.
@@ I-A Basic Notation
Table I: Modeling notation
I-B Preliminaries
In the following, we introduce relevant concepts from distributionally robust optimization (DRO) [27]. The conditional value at risk (CVaR) is a popular risk measure that provides the average of the tail end of the loss distribution. Intuitively, it represents the expected average of the worst-case outcomes above a certain threshold $\beta\in[0,1]$. It is defined as follows.
@@ II Problem setting
Crucially, the aggregate load profile of each class $c$ is unknown at the beginning of the planning horizon. Hence, we treat it as a random vector $s_{c}$ drawn from an underlying probability distribution, denoted by $\mathbb{P}$. Each element $s_{c,k}$, with $k\in\mathcal{H}$, within this random vector represents the aggregate load of compute jobs of class $c$ submitted at time $k$ for processing.
@@
$\sum_{d\in\mathcal{D}}\sum_{t\in\mathcal{T}}Y_{k,c,t,d}\geq 1,\quad\forall k\in\mathcal{H},\;c\in\mathcal{C}.$ (3)
@@
$Y_{k,c,t,d}=0,\quad\forall k\in\mathcal{H},c\in\mathcal{C},t\in\mathbb{Z}_{[k:k+h_{c}]},d\in\mathcal{D}.$ (4b)
Figure 1: An example of the aggregate load schedule for compute loads from two flexibility classes $c\in\{1,2\}$ at a single cluster $d=1$ over $24$ hours. The VCC (solid line) $v_{t,1}$ limits the allocable load at each time interval $t$, while the true capacity $\overline{v}_{t,d}$ (dashed line) is obtained by subtracting the inflexible load2 from the (cluster) machine capacity.
The operational capacities $v_{t,d}$ of each cluster $d$, known as virtual capacity curve (VCC), determine the largest load allocable at a cluster. Inherently, each $v_{t,d}$ is limited by the true capacity $\overline{v}_{t,d}$, which takes into account the (cluster) machine capacity and the predicted inflexible load [^2] Inflexible load consists of compute jobs that cannot be shifted in time and across-locations. yielding the following constraint
@@
In the current Google model [26], the VCCs were computed independently of the job scheduling strategy. In this paper, we treat $v=\text{col}(v_{t,d})_{d\in\mathcal{D},t\in\mathcal{T}}$ as an additional decision variable and propose the co-design of VCCs and scheduling strategy $Y$. Intuitively, the allocated aggregate load per cluster $d$ and time $t$ must be lower than the VCC limit $v$. Since the compute loads $s_{k,c}$ are random variables, we enforce this constraint in a probabilistic sense, using the following Conditional Value-at-Risk (CVaR) constraint:
$\text{CVaR}_{1-\beta}^{s\sim\mathbb{P}}\left[\max_{\begin{subarray}{c}t\in\mathcal{T},\\ d\in\mathcal{D}\end{subarray}}\left(\sum_{c\in\mathcal{C}}\sum_{k\in\mathcal{H}}Y_{k,c,t,d}\cdot s_{k,c}-v_{t,d}\right)\right]\leq 0.$ (6)
@@
The objective is a piece-wise linear function of compute usage and can incorporate different time-varying grid signals (carbon intensities, carbon free energy score, electricity prices, DR signals, etc.) as well as infrastructure related costs (peak power, stand-by machine cost etc.). Here, we use a similar objective function to Google’s real implementation in [26], considering carbon footprint and daily peak power as a function of VCCs:
$f(v)=\sum_{t\in\mathcal{T}}\sum_{d\in\mathcal{D}}\rho^{\text{carb}}_{t,d}\leavevmode\nobreak\ v_{t,d}+\sum_{d\in\mathcal{D}}\rho^{\text{in}}_{d}\,\|\text{col}(v_{t,d})_{t\in\mathcal{T}}\|_{\infty},$ (7)
@@
where $\rho^{\text{carb}}_{t,d}$ is a metric for the generated carbon footprint [30] and $\rho_{d}^{\text{in}}$ is associated with infrastructure costs driven by a cluster’s peak power consumption. The carbon impact of utilizing computing power is modeled using a linear relationship, as described in [16, 21]. As predictions for time-varying grid signals can be inaccurate or change rapidly, so re-optimizing throughout the day is necessary for better performance and to participate in DR schemes.
@@
where $y=\text{col}(y_{r})_{r=0}^{KCTD}$, $y_{r}=Y_{k,c,t,d}$ with $r=K(c-1)+k+KC(t-1)+KCT(d-1)$ is a vectorized version of the load schedule, $Y$, which is more suitable as an optimization variable. The constraint (8b) is a reformulation of (6) in terms of $y$ with $F$ defined as
@@
where the matrix $A_{td}\in\mathbb{R}^{KCTD\times KC}$ has ones in its $(kctd,kc)$ entries and zeroes elsewhere, while the vector $b_{td}\in\mathbb{R}^{TD}$ has its $td$ entries set to $-1$ and zero elsewhere. An example of a load schedule is shown in Figure 1.
@@
The stochastic problem in (8a) could be formulated only in terms of scheduling strategy $y$ and the random load profile $s$. The epigraph reformulation we chose through constraint (6) gives an explicit CVaR bound $v_{t,d}$ for each cluster $d$ and time $t$ for the random aggregate load profile $s_{k,c}$. As done by Google [26], the virtual capacities $v_{t,d}$ can be enforced as hard constraints on the executable load in each cluster. $\square$
@@
III Distributionally-robust scheduling

### v3

@@ Abstract
Data centers are significant contributors to carbon emissions and can strain power systems due to their high electricity consumption. To mitigate this impact and to participate in demand response programs, cloud computing companies strive to balance and optimize operations across their global fleets by making strategic decisions about when and where to place compute jobs for execution. In this paper, we introduce a load shaping scheme which reacts to time-varying grid signals by leveraging both temporal and spatial flexibility of compute jobs to provide risk-aware management guidelines and job placement with provable performance guarantees based on distributionally robust optimization. Our approach divides the problem into two key components: (i) day-ahead planning, which generates an optimal scheduling strategy based on historical load data, and (ii) real-time job placement and (time) scheduling, which dynamically tracks the optimal strategy generated in (i). We validate our method in simulation using normalized load profiles from randomly selected Google clusters, incorporating time-varying grid signals. We can demonstrate significant reductions in carbon cost and peak power with our approach compared to myopic greedy policies, while maintaining computational efficiency and abiding to system and grid constraints.
@@ I Introduction
The literature covers various aspects of the problem but has several key limitations. Specifically, most works (i) do not exploit temporal and spatial flexibility jointly [19, 20, 21, 22]; (ii) rely on either optimization-based planning approaches, which are not tractable in real-time operation [21, 23], or myopic greedy policies, which are suboptimal as they do not take into account future grid signals or compute load predictions [24, 16, 25]; and (iii) either assume exact knowledge of incoming compute jobs, which is unrealistic, or lack specifics on real-time job placement [21, 18].
@@
Most real-time job-placement schemes rely on lightweight but inherently myopic greedy rules. For example, [26] chooses, at every job arrival, the cluster that minimizes a weighted score of instantaneous renewable supply, electricity price, power-of-usage effectiveness, and server fragmentation, achieving improved energy-cost savings without any look-ahead. The work in [27] maximizes on-site solar utilization by deferring deadline-flexible batch jobs to daylight hours. The authors in [28] push this idea further: they combine dynamic electricity prices with minute-ahead PV “now‑casting” and route workloads among geographically separated data centers that are expected to have the cheapest (or greenest) power supply in the near future. The authors in [29] apply an $O(1)$ rule that directs each request to the site minimizing an energy-delay product, thereby trading off incremental energy use against network-induced latency. Although all four heuristics are computationally cheap, they optimize each placement in isolation and provide no formal guarantees on efficiency or robustness under stochastic compute-load conditions.
@@
We propose a distributionally-robust day-ahead planning problem coupled with a real-time placement algorithm which is driven by data, offers robustness guarantees, and considers various degrees of flexibility of jobs while explicitly incorporating grid codes and demand response (DR) events. Integrating stochasticity of DR events is very relevant as Google has been called in the past to drop load in emergency situations [15]. To summarize, our approach is unique in the following ways:
@@
    Data-driven: We directly approximate the probability distribution of compute loads from data within the planning problem, without making assumptions on the distribution.
@@
  5. 5.
    Grid requirements: We incorporate grid codes and demand response services in the modeling of the scheduling problem in a probabilistic sense, to account for the unknown distribution of demand response events and compute loads.
@@
Although this work focuses on real-time operation of data centers, the ability to reliably quantify and manage flexible compute demand has direct implications also for long-term planning in power systems. For example, when incorporated into generation–transmission expansion models, the spatio-temporal flexibility of data centers can (i) cut peak demand and thus defer or reduce new generation capacity [31]; (ii) increase the use of spatially diverse wind and solar, informing renewable siting and curtailment mitigation [32]; and (iii) enable co‑optimized expansion models in which demand flexibility lowers total system costs [33]. Hence, the operational approach developed here provides not only short-term carbon and cost benefits but also actionable inputs for the long-term integrated planning of grid infrastructure and compute facilities.
The rest of the paper is structured as follows. We introduce basic notation and state preliminaries in Section II. In Section III, we introduce the model and formulate the stochastic optimization problem. In Section IV, we derive a distributionally-robust load schedule, approximating the ambiguity set from historical data. Section V discusses real-time job placement. Section VI and Section VII present simulation results and Section VIII concludes the paper.
II Notation and Preliminaries
We denote the set of the first $K$ positive integers as $\mathbb{Z}_{K}:=\{1,\dots,K\}$, and the set of positive integers from $t$ to $T$, with $t<T$, as $\mathbb{Z}_{[t:T]}:=\{t,\dots,T\}$. The operator $\cdot\,]_{+}:=\max\{\cdot,0\}$ is the projection on the positive orthant. Given some variables $v_{t,d}$, with $t\in\mathcal{T}$ and $d\in\mathcal{D}$, we denote the stacked vector as $v:=\text{col}(v_{t,d})_{t\in\mathcal{T},d\in\mathcal{D}}=[v_{1,1},\dots,v_{T,D}]^{\top}$. Table I summarizes all parameters, sets, and variables used.
Table I: Modeling notation
@@ II Notation and Preliminaries
$p_{t,d}$ Power load
$\bar{p}_{d}$ Maximum power limit
$\bar{p}_{d}^{\uparrow/\downarrow}$ Maximum ramp up/down power limit
$\Delta t$ Time conversion factor from ramping to planning time scales
$\alpha_{1,d}$, $\alpha_{0,d}$ Slope and intercept of the affine power model
@@
In the following, we introduce relevant concepts from distributionally robust optimization (DRO) [34]. The conditional value at risk (CVaR) is a popular risk measure that provides the average of the tail end of the loss distribution. Intuitively, it captures the expected value of the worst-case $\beta$ fraction of outcomes with $\beta\in(0,1]$. It is defined as follows.
@@ III Problem setting
Crucially, the exact future aggregate load profile of each class $c$ is unknown at the beginning of the planning horizon. Hence, we treat it as a random vector $s_{c}$ drawn from an underlying probability distribution, denoted by $\mathbb{P}$. Each element $s_{c,k}$, with $k\in\mathcal{H}$, within this random vector represents the aggregate load of compute jobs of class $c$ submitted at time $k$ for processing.
@@ III-A Compute load constraints
$\sum_{d\in\mathcal{D}}\sum_{t\in\mathcal{T}}Y_{k,c,t,d}\,=\,1,\quad\forall k\in\mathcal{H},\;c\in\mathcal{C}.$ (3)
@@
$Y_{k,c,t,d}=0,\quad\forall k\in\mathcal{H},c\in\mathcal{C},t\in\mathcal{T}\backslash\mathbb{Z}_{[k:k+h_{c}]},d\in\mathcal{D}.$ (4b)
Figure 1: An example of the aggregate load schedule for compute loads from two flexibility classes $c\in\{1,2\}$ at a single cluster $d=1$ over 2424 hours. The VCC (solid line) $v_{t,1}$ limits the allocable load at each time interval $t$, while the true capacity $\overline{v}_{t,d}$ (dashed line) is obtained by subtracting the inflexible load2 from the (cluster) machine capacity.
The operational capacities $v_{t,d}$ of each cluster $d$, known as virtual capacity curve (VCC), determine the largest load allocable at a cluster. The VCC limits induce indirect temporal shifting of compute jobs as low capacity during times of peak compute load will lead to jobs being queued and their execution being delayed.
Inherently, each $v_{t,d}$ is limited by the true capacity $\overline{v}_{t,d}$, which takes into account the (cluster) machine capacity and the predicted inflexible load [^2] Inflexible load consists of compute jobs that cannot be shifted in time and across-locations. yielding the following constraint
@@
In the current Google model [30], the VCCs were computed independently of the job scheduling strategy. In this paper, we treat $v=\text{col}(v_{t,d})_{d\in\mathcal{D},t\in\mathcal{T}}$ as an additional decision variable and propose the co-design of VCCs and scheduling strategy $Y$. Intuitively, the allocated aggregate load per cluster $d$ and time $t$ must be lower than the VCC limit $v$, thus we define the following constraint function
$g_{1}(Y,s,v)=\max_{\begin{subarray}{c}t\in\mathcal{T},\\ d\in\mathcal{D}\end{subarray}}\left(\sum_{c\in\mathcal{C}}\sum_{k\in\mathcal{H}}Y_{k,c,t,d}\cdot s_{k,c}-v_{t,d}\right)$ (6)
Since the compute loads $s_{k,c}$ are random variables, we enforce this constraint in a probabilistic sense as following using the following Conditional Value-at-Risk (CVaR) constraint:
$\text{CVaR}_{1-\beta}^{s\sim\mathbb{P}}\left[\max_{\begin{subarray}{c}t\in\mathcal{T},\\ d\in\mathcal{D}\end{subarray}}g_{1}(Y,s,v)\right]\leq 0.$ (7)
@@
III-B Power constraints
o effectively manage power consumption of a cluster, it is necessary to model the relationship between compute load and power. We will make use of the power model that is running in production and relates CPU usage of Google’s Power Distribution Units (PDU) to their power consumption [37]. A single PDU is comprised of thousands of machines, and a handful of PDUs comprise a cluster. Although DC infrastructure planning and real-time job placement depend on multi-dimensional resource usage (CPU, RAM, disk usage), it was demonstrated in [37] that an affine model applies for PDUs with different machine types (e.g. standard compute, storage, networking). Therefore, we can write
$p^{\text{PDU}}=\alpha_{1}^{r}\,l^{\text{PDU}}+\alpha_{0}^{r},$
where $l^{\text{PDU}}$ is the CPU usage at a PDU and the slope $\alpha_{1}^{r}$ and intercept $\alpha_{0}^{r}$ would be fitted for each PDU and usage regimes. The idle power of machines depends on the platform family and lies between 30%-40% of their maximum power consumption, i.e., $\alpha_{0}\in[0.3,0.4]\cdot p_{\text{max}}^{\text{PDU}}$. The slope is primarily influenced by the CPU usage and the workload mix running in the PDU, and it is in most cases smaller than 10 (most commonly smaller than 1), thus $\alpha_{1}^{r}\in[1,10]$. However, in this work, we are interested in the power consumption of an entire cluster at time $t$, denoted as $p_{t,d}$, which is the aggregate power consumption of all PDUs in a cluster. In practice, DCs are connected to the power grid and must adhere to grid codes, such as contractual maximum power consumption limits $\bar{p}_{d}$, to avoid getting their supply throttled. This yields the constraint
$0\leq\text{PUE}\cdot\,p_{t,d}\;\leq\;\bar{p}_{d},\quad\forall t\in\mathcal{T},\,d\in\mathcal{D}.$ (8)
where PUE is the Power Usage Efficiency factor which considers power required for cooling of the data center and translates the aggregate power consumption of PDUs into total grid-level power demand. In practice, Google maintains a nearly constant PUE of 1.09 across all data centers and seasons [38].
Remark 2.
Google data centers rely on local controllers for cooling [39] which achieve constant PUE levels across seasons and locations [38]. Therefore, it is not necessary to explicitly model cooling power in our load shifting mechanism and incorporating the proportional PUE gain into (8) suffices.
Another relevant grid constraint for DCs are ramping limits which specify how fast load may be dropped or increased without destabilizing the local grid. For instance the Swiss code specifies a linear change over a period of 10 minutes [40]. The exact time scales and limits specified in the code differ across grids. Yet, most of them can be modeled using a linear constraint as follows
$\Delta t\,\bar{p}_{d}^{\downarrow}\leq(p_{t+1,d}-p_{t,d})\;\leq\;\Delta t\,\bar{p}_{d}^{\uparrow},\quad\forall t\in\mathcal{T},\,d\in\mathcal{D},$ (9)
where $\bar{p}_{d}^{\uparrow/\downarrow}$ are ramping up and down limits and $\Delta t$ is a time conversion factor between ramping and planning time scales. This approach is consistent with ramping limits considered in the literature [18] [^3] Note that multiple clusters comprise one data center, and thus depending on the infrastructure set-up, DC limits would have to be converted to per-cluster limits.. Similarly to the VCC limits, the power consumed at a given cluster is determined by the random, unknown load profile $s_{k,c}$ and the scheduling strategy $Y$. To enforce this as a probabilistic constraint, we define the function
$g_{2}(Y,s,p)=\max_{\begin{subarray}{c}t\in\mathcal{T},\\ d\in\mathcal{D}\end{subarray}}\left(\alpha_{1,d}\sum_{c\in\mathcal{C}}\sum_{k\in\mathcal{H}}Y_{k,c,t,d}\cdot s_{k,c}+\alpha_{0,d}-p_{t,d}\right),$ (10)
where $\alpha_{1,d}$ is the slope and $\alpha_{0,d}$ the idle power consumption of an entire cluster $d$. Having both constraint functions $g_{1}$ defined in (6) and $g_{2}$ in place enables us to combine them in one probabilistic constraint using a single CVaR measure as follows:
$\text{CVaR}_{1-\beta}^{s\sim\mathbb{P}}\left[\max_{\begin{subarray}{c}t\in\mathcal{T},\\ d\in\mathcal{D}\end{subarray}}\left(g_{1}(Y,s,v),g_{2}(Y,s,p)\right)\right]\leq 0.$ (11)
Note that this formulation constrains the worst case violation in terms of VCC limits and power constraints jointly and avoids introducing additional conservatism which would be the case if one added another independent CVaR constraint.
III-C Objective function
@@
The objective is a piece-wise linear function of the VCC limits and can incorporate different time-varying grid signals (carbon intensities, carbon free energy score, electricity prices, DR signals, etc.) as well as infrastructure related costs (peak power, stand-by machine cost etc.). We provide a summary of different cost components which are relevant for data center load shifting and have been used in the literature in Table II. Here, we use a similar objective function to Google’s real implementation in [30], considering carbon footprint and daily peak power as a function of VCCs:
$f(v)=\sum_{t\in\mathcal{T}}\sum_{d\in\mathcal{D}}\rho^{\text{carb}}_{t,d}\penalty 10000\ v_{t,d}+\sum_{d\in\mathcal{D}}\rho^{\text{in}}_{d}\,\|\text{col}(v_{t,d})_{t\in\mathcal{T}}\|_{\infty},$ (12)
@@
where $\rho^{\text{carb}}_{t,d}$ is a metric for the generated carbon footprint [41] and $\rho_{d}^{\text{in}}$ is associated with infrastructure costs driven by a cluster’s peak power consumption. Peak power is used as an important signal to guide when to increase capacity by adding more machines in a cluster. The carbon impact of utilizing computing power is modeled using a linear relationship, as described in [16, 21]. As predictions for time-varying grid signals can be inaccurate or change rapidly, so re-optimizing throughout the day is necessary for better performance and to participate in DR schemes. The parameter $\rho^{\text{carb}}_{t,d}$ ideally represents how 1 MW increase in load at time $t$ and cluster $d$ affects the system’s total carbon emissions. However, it is important to note that carbon intensity data is generally not available in real time. Instead, estimates must be made using the absolute values provided by [41].
Table II: Metrics for load‑shifting decisions in data centers.
Description Publication
Generator cost [42]
Dispatch/power cost [18, 26]
Carbon intensity (local, average, marginal) [23, 43]
Migration of load (delay, network charges) [23, 43, 44]
Fragmentation cost [42]
Max power or peak load [23]
III-D Demand response
Google is offering demand response (DR) services in DCs across the globe [15]. During extreme weather events in Nebraska Google has already successfully supported the local utility in avoiding a blackout by reducing power consumption [15]. Beyond monetary incentives, DC operators may need to offer flexibility to utilities in regions with strained power grids if they wish to secure building permissions and sufficient power capacity in the future. Typically, the local distribution system operator will provide a demand-response signal day-ahead, and Google would declare compute capacity in MW that can be dropped if requested. Thus, the price signal is available in advance but if and when the demand response event happens during the next day is uncertain. We model such events using scenarios $s_{k,\text{DR}}$ of a new class without spatial or temporal flexibility. Thus, any DR request will be placed immediately in the given cluster. All constraints (3)-(11) remain unchanged, but now apply to the extended class set $\tilde{\mathcal{C}}:=\{1,\dots,C,\text{DR}\}$, which includes the additional DR class. The advantage of treating DR events using a stochastic model is that the job completion requirements still hold in expectation as defined by the CVaR constraint.
Remark 3.
Without available data, capacity for DR events must be reserved across all time steps, effectively leading to a uniform robustification. However, when historical data or prior assumptions about DR events is accessible, capacity can be strategically reserved only for periods with a high probability of a DR event occurring, thereby reducing conservatism and improving cost performance.
III-E Stochastic scheduling problem
@@
where $y=\text{col}(y_{r})_{r=0}^{KCTD}$, $y_{r}=Y_{k,c,t,d}$ with r=K(c1)+k+KC(t1)+KCT(d1)r=K(c-1)+k+KC(t-1)+KCT(d-1) is a vectorized version of the load schedule, $Y$, which is more suitable as an optimization variable. The constraint (13b) is a reformulation of (7) in terms of $y$ with $F$ defined as
@@
where the matrix $A_{td}\in\mathbb{R}^{KCTD\times KC}$ has ones in its $(kctd,kc)$ entries and zeroes elsewhere, while the vector $b_{td}\in\mathbb{R}^{TD}$ has its $td$ entries set to 1-1 and zero elsewhere. An example of a load schedule is shown in Figure 1. Note that we omit adding the constraints for the power limit (8) and ramping limit (9) in the scheduling problem (13) and the subsequent theoretical derivations for improved clarity. However, the derivations with constraints (8)-(11) included would follow the same structure.
@@
The stochastic problem in (13a) could be formulated only in terms of scheduling strategy $y$ and the random load profile $s$. The epigraph reformulation we chose through constraint (7) gives an explicit CVaR bound $v_{t,d}$ for each cluster $d$ and time $t$ for the random aggregate load profile $s_{k,c}$. As done by Google [30], the virtual capacities $v_{t,d}$ can be enforced as hard constraints on each cluster’s executable load in each cluster. $\square$
@@
IV Distributionally-robust scheduling